SJISCONV(1) SJISCONV(1) NAME sjisconv−convertaTeXdocumentinSJISencodinginto‘preprocessed’form. SYNOPSIS sjisconv<infile>outfile DESCRIPTION SJISencodingforJapaneseusesthecharacters{,},and\whichhavespecialmeaningsinTEX documents. AfterprocessingaLATEX2εdocumentwhichcontainsSJIScharacterswithsjisconvyouneednot careaboutthesespecialcharacters. ThisfilterispartoftheCJKmacropackageforLATEX2ε. SEEALSO bg5conv(1),cefconv(1),cef5conv(1),cefsconv(1),extconv(1), theCJKdocumentationfiles. AUTHOR WernerLemberg<wl@gnu.org> CJKVersion4.3.0 20-Jun-1999 1